# Entanglement in the Schwinger effect

### v1

@@ Abstract
We analyze entanglement generated by the Schwinger effect using a mode-by-mode formalism for scalar and spinor QED in constant backgrounds. Starting from thermal initial states, we derive compact, closed-form results for bipartite entanglement between particle–antiparticle partners in terms of the Bogoliubov coefficients. For bosons, thermal fluctuations enhance production but suppress quantum correlations: the logarithmic negativity is nonzero only below a (mode-dependent) critical temperature $T_{c}$. At fixed $T$, entanglement appears only above a critical field $E_{\text{crit,entang}}$. For fermions, we observe a qualitatively different pattern: at finite $T$ entanglement exists only within a finite window $E_{\text{min}}<E<E_{\text{max}}$, with a temperature-independent optimal field strength $E_{*}$ that maximizes the logarithmic negativity. Entanglement is vanishing above $T_{\text{max}}=\omega/\text{arcsinh}(1)$. We give quantitative estimates for analog experiments, where our entanglement criteria convert directly into concrete temperature and electric field constraints. These findings identify realistic regimes where the quantum character of Schwinger physics may be tested in the laboratory.
@@ I Introduction
The Schwinger effect – the spontaneous creation of particle-antiparticle pairs from the vacuum in the presence of a strong electric field – is one of the most distinctive predictions of quantum field theory under extreme conditions. Originally formulated in the context of quantum electrodynamics (QED) [Schwinger:1951nm], it shares deep connections with other interesting phenomena such as cosmological particle production [Parker:2025jef, Parker:1968mv, Parker:1969au] and the Hawking effect [Hawking:1974rv, Hawking:1975vcx]. These processes not only signal an instability of the vacuum, but also imprint non-trivial quantum correlations in the created particles, which can be used, in principle, to probe the quantum nature of the underlying dynamics.
On the experimental side, there is growing interest in experimentally probing non-perturbative effects in QED. High-intensity laser facilities aim to reach the critical field strength where vacuum pair production becomes observable [Dunne:2008kc, Karbstein:2019oej, Fedotov:2022ely]. At the same time, analogue experiments have emerged as powerful tools to simulate aspects of strong-field QED in condensed matter systems. The mesoscopic Schwinger effect provides an excellent playground to study properties of the (analogue) Schwinger effect for fermions, since a universal Schwinger conductance in graphene was reported [Schmitt:2022pkd]. For bosons, magnetic inhomogeneities in chiral magnets have been proposed as a route to Schwinger-like magnon production [Hongo:2020xaw, Adorno:2023olb].
@@
Building on this framework, we show that the particle-antiparticle pairs generated via vacuum polarization necessarily carry entanglement that cannot be reproduced by classical stochastic processes or thermal noise. Moreover, we quantify how much this entanglement degrades in the presence of a thermal background, identifying a critical temperature above which the entanglement vanishes, even though pair production persists. This provides an operational criterion to distinguish the quantum nature of the process.
@@
At a fixed temperature, we also analyze entanglement in terms of the electric field. For bosons, there is only entanglement if the electric field is above a critical value $E_{\text{entang}}$. For fermions, the pattern is different: pairs are entangled only if the electric field is within a finite window $E_{\text{min}}<E<E_{\text{max}}$. Furthermore, we also find an optimal electric field for entanglement $E_{*}$, which is temperature independent. We also study a mechanism – valid for scalar fields – to enhance entanglement by tuning the initial state. Specifically, we show that by preparing the vacuum in a squeezed state, one can increase the quantum entanglement produced during the pair creation process of bosons. This provides a way to control quantum correlations in strong-field pair production, with potential applications in analogue experiments.
@@
The entanglement structure of pairs produced via the Schwinger mechanism has been previously studied in various settings, using different entanglement measures such as the von Neumann entropy, and logarithmic negativity. Previous works have focused on both scalar and spinor fields in constant or pulsed backgrounds at zero temperature [Ebadi2014, Li_2017, Bhattacharya2020, Kaushal_2025]. In this work, we study in full detail the interplay between different input states, such as thermal and single-mode squeezed states, investigating how they affect the entanglement generated in the Schwinger effect, thus providing a holistic analysis of the Schwinger process. Beyond strong QED, recent efforts in collider physics [Afik:2025ejh, Aguilar-Saavedra:2022uye, Aguilar-Saavedra:2022wam] and early-universe cosmology [Pueyo:2024twm, Bhardwaj:2023squ] have explored how entanglement-based observables can be used to test quantum correlations in fundamental processes. Similar ideas have been recently discussed in the context of gravitational radiation [Manikandan:2025qgv, Manikandan:2025hlz].
@@
The outline of the article is as follows. In Section II we briefly review the Schwinger effect for both scalars and fermions following a mode-by-mode approach in terms of the Bogoliubov coefficients. This approach is particularly useful for understanding quantum entanglement. We focus on two cases: i) constant electric field $\vec{E}$, and ii) constant and anti-parallel electric and magnetic fields $\vec{E}$ and $\vec{B}$. We then introduce the tools needed to quantify quantum entanglement, including the Gaussian formalism, and the definition of logarithmic negativity as our main measure of entanglement in Section III. In Section IV we present our main results: a comparative study of scalar and spinor entanglement in QED, the impact of thermal initial states, and the identification of critical temperatures and electric field strengths above (or under) which entanglement vanishes. We also discuss how these findings can be used in future experiments. Finally we summarize our conclusions in Section V.
@@ II Schwinger effect
The instability of the quantum vacuum in strong electric fields was first anticipated by Heisenberg and Euler through an effective action approach for constant backgrounds [Heisenberg:1936nmg] (see Ref. [Weisskopf:1936hya] for scalar QED). Schwinger later formalized this result using the modern language of QED by computing the imaginary part of the one-loop effective action to evaluate the vacuum-persistence amplitude [Schwinger:1951nm],
@@
with $V\Delta t$ the spacetime volume, and $e$, $m$, $E$ the particle charge, mass, and the external field strength, respectively. This expression reveals the exponential suppression of the process for subcritical fields, and defines the natural scale $E_{\mathrm{crit}}=\pi m^{2}/e$ as the threshold where non-perturbative pair production becomes significant (see Ref. [Dunne:2004nc] for an historical perspective and Ref. [Schubert:2024vft] for a review on both historical and technical aspects).
In what follows we adopt the (equivalent) mode–function viewpoint based on frequency mixing and Bogoliubov transformations. [^2] See Refs. [Lozanov:2018kpk, Domcke:2019qmm, Ferreiro:2018qdi, Beltran-Palau:2020hdr, Alvarez-Dominguez:2023ten, Alvarez-Dominguez:2023zsk] for different aspects of the Schwinger effect within this formalism. And Ref. [VicenteGarcia-Consuegra:2025lkh] for a stochastic generalization in the effective action formalism. By analyzing the frequency mixing of solutions to the Klein–Gordon and Dirac equations in time-dependent backgrounds, we can identify the natural in and out mode bases and compute the coefficients that relate them. This provides direct access to the particle content of the theory and, crucially for our purposes, to the quantum correlations between the produced pairs.
@@ II.1 Bosons
The appropriate inner product in the space of solutions between two field configurations $\phi_{1}$ and $\phi_{2}$ is given by the conserved symplectic form [^4] The complex-conjugate field carries opposite charge and its (natural) KG product uses $D_{\mu}^{*}=\partial_{\mu}-ieA_{\mu}$. As a result, if $\phi$ has positive KG norm, $\phi^{*}$ has negative KG norm with respect to the conjugate product (see Ref. [Alvarez-Dominguez:2025nce] for further details).
@@ II.2 Fermions
Unlike the bosonic case, and for the sake of brevity, we do not treat here the $\vec{B}=0$ configuration separately, since the Bogoliubov coefficients can be obtained from the case with anti-parallel $\vec{E}$ and $\vec{B}$ fields in the limit $B\to 0$. More details can be found in Appendix A. We proceed by decomposing the spinor field into Fourier modes along the $y$ and $z$ directions, and expanding its $x$-dependence in terms of Hermite polynomials: [^5] The parameter $l$ of the Hermite polynomials is often called “Landau level” (for instance in [Domcke:2019qmm]).
@@ II.3 Bogoliubov coefficients
are the well-known Bogoliubov coefficients, that can be computed as inner products of the in and out solutions[parker-toms]. [^6] See Ref. [Padmanabhan:1991uk] for a detailed computation of the Bogoliubov coefficients in the constant $E$-background case, and for interesting remarks on different gauge choices, the relationship with cosmological particle production, and connection to effective action methods. Using the fact that $|\Gamma(\frac{1}{2}+\frac{i\lambda}{2})|^{2}=\frac{\pi}{\cosh(\frac{\lambda\pi}{2})}e^{\lambda\pi/2}$ one can verify that
@@
Let us move now to the spinor case. Just as in the scalar case, the relationship between the in the out basis comes from the relationship between the different $S_{i}$ functions. In the $\vec{B}\parallel\vec{E}$ case, we find the following Bogoliubov transformations between the creation and annihilation operators at early and late times (in agreement with Refs. [Nikishov:1969tt, Soldati:2011gi])
@@
where the Bogoliubov coefficients read (see, for example, Ref. [Soldati:2011gi] for a detailed derivation)
@@ III Quantum entanglement
As we are interested in quantifying the entanglement in Schwinger effect, both for bosons and fermions, we provide a pedagogical presentation of the tools and formalism we leverage from quantum information theory. We begin by defining some of the most popular entanglement and correlations quantifiers we use in our calculations, more specifically, entanglement entropy, logarithmic negativity, and mutual information, and then we illustrate how to compute the entanglement in some specific cases of interest.
@@ III.1 Quantifying entanglement
Quantifying the amount of entanglement in quantum states is a topic of central interest in quantum information theory and there exists a whole program of entanglement quantification with several measures [Plenio:2006] being employed for this task. The most widely known, among them, is the von Neumann entropy of the reduced state, computed by $S_{V}=-\text{Tr}_{A}[\hat{\rho}\log_{2}\hat{\rho}$], where $\text{Tr}_{A}$ denotes the partial trace with respect to subsystem $A$ (the result is independent on whether tracing out subsystem $A$ or $B$). This measure, however, is limited to the cases where the total state $\hat{\rho}$ is pure. This is rarely the case in real-life setups since pure states are associated to isolated systems. In realistic frameworks, systems are always exposed, to some degree, to noise resulting from interactions between the system of interest and its environment, i.e. the set of undesired degrees of freedom.
@@ III.2 Gaussian formalism of continuous variable systems
Continuous variable systems are those whose observables possess a continuous spectrum, such as the harmonic oscillator. Upon quantization, the states of such systems are elements of an infinite-dimensional Hilbert space, and, hence, computing entanglement in such systems amounts to manipulating infinite-dimensional density matrices. Gaussian systems, however, such as those generated by quadratic Hamiltonians, enjoy a drastic simplification: they are fully characterized by their first and second moments, whose size depends on the number of degrees of freedom. We now briefly review the Gaussian formalism and how to employ it to compute entanglement. Given that we are interested in the mode of pairwise entanglement, we will restrict the discussion to the case of two degrees of freedom. A pedagogical reference for the Gaussian state formalism and entanglement in Gaussian states can be found in [Serafini:book].
@@
Now, let us turn our attention to the logarithmic negativity. Simon extended the PPT criterion to two-mode continuous variable Gaussian systems [Simon:2000]. The key step towards computing entanglement is the operation of the partial transposition. In terms of a $4\times 4$ covariance matrix $\bm{\sigma}$, the partial transposition reads [Serafini:book]
@@ III.3 Example: a two-level bipartite system
$\hat{\rho}^{(\text{out})}=\begin{pmatrix}|\alpha|^{2}&0&0&\alpha^{*}\beta^{*}\\ 0&0&0&0\\ 0&0&0&0\\ \alpha\beta&0&0&|\beta|^{2}\end{pmatrix}\,$. (47)
@@
where $\omega_{i}$ is the energy of the $i$-th particle and $f_{i}=\left(e^{\omega_{i}/T}+1\right)^{-1}$ is the mean occupation number of the $i$-th particle [^10] We use $f$ to emphasize that they have Fermi-Dirac statistics. We consider $f_{1}=f_{2}=f$. The total state reads
@@
Following a similar procedure as for the vacuum case, one finds that the final covariance matrix reads

### v2

@@ Abstract
We analyze entanglement generated by the Schwinger effect using a mode-by-mode formalism for scalar and spinor QED in constant backgrounds. Starting from thermal initial states, we derive compact, closed-form results for bipartite entanglement between particle–antiparticle partners in terms of the Bogoliubov coefficients. For bosons, thermal fluctuations enhance production but suppress quantum correlations: the logarithmic negativity is nonzero only below a (mode-dependent) critical temperature $T_{c}$. At fixed $T$, entanglement appears only above a critical field $E_{\text{entang}}$. For fermions, we observe a qualitatively different pattern: the fermionic logarithmic negativity is non-vanishing at finite temperature, and is monotonically suppressed by thermal noise. As a function of the electric field, it is non-monotonic, featuring a temperature-independent optimal field strength $E_{*}$ and decreasing on both sides of the maximum. We give quantitative estimates for analog experiments, where our entanglement criteria convert directly into concrete temperature and electric field constraints. These findings identify realistic regimes where the quantum character of Schwinger physics may be tested in the laboratory.
preprint: TUM-HEP-1582/25
@@ I Introduction
The Schwinger effect – the spontaneous creation of particle-antiparticle pairs from the vacuum in the presence of a strong electric field – is one of the most distinctive predictions of quantum field theory under extreme conditions. Originally formulated in the context of quantum electrodynamics (QED) [68], it shares deep connections with other interesting phenomena such as cosmological particle production [57, 55, 58] and the Hawking effect [38, 39]. These processes not only signal an instability of the vacuum, but also imprint non-trivial quantum correlations in the created particles, which can be used, in principle, to probe the quantum nature of the underlying dynamics.
On the experimental side, there is growing interest in experimentally probing non-perturbative effects in QED. High-intensity laser facilities aim to reach the critical field strength where vacuum pair production becomes observable [24, 45, 26]. At the same time, analogue experiments have emerged as powerful tools to simulate aspects of strong-field QED in condensed matter systems. The mesoscopic Schwinger effect provides an excellent playground to study properties of the (analogue) Schwinger effect for fermions, since a universal Schwinger conductance in graphene was reported [65]. For bosons, magnetic inhomogeneities in chiral magnets have been proposed as a route to Schwinger-like magnon production [42, 2]. [^1] For complementary perspectives on entanglement in Schwinger pair creation in holographic setups, see [35, 34], building on the seminal proposals [44, 74].
@@
Building on this framework, we show that the particle-antiparticle pairs generated via vacuum polarization necessarily carry entanglement that cannot be reproduced by classical stochastic processes or thermal noise. Moreover, we quantify how much this entanglement degrades in the presence of a thermal background, identifying a critical temperature above which the (bosonic) logarithmic negativity vanishes, even though pair production persists. This provides an operational criterion to distinguish the quantum nature of the process. For fermions, we employ the fermionic logarithmic negativity proposed in Refs. [71, 70] and find that entanglement is always present at finite temperature, although it is strongly suppressed as thermal noise increases.
@@
At a fixed temperature, we also analyze entanglement in terms of the electric field. For bosons, there is only entanglement if the electric field is above a critical value $E_{\text{entang}}$. For fermions, the pattern is different: the logarithmic negativity is a non-monotonic function of the electric field, featuring a temperature-independent optimal field strength $E_{*}$, and decays for bigger and smaller values of $E$. We also study a mechanism – valid for scalar fields – to enhance entanglement by tuning the initial state. Specifically, we show that by preparing the vacuum in a squeezed state, one can increase the quantum entanglement produced during the pair creation process of bosons. This provides a way to control quantum correlations in strong-field pair production, with potential applications in analogue experiments.
@@
The entanglement structure of pairs produced via the Schwinger mechanism has been previously studied in various settings, using different entanglement measures such as the von Neumann entropy, and bosonic logarithmic negativity. Previous works have focused on both scalar and spinor fields in constant or pulsed backgrounds at zero temperature [25, 49, 15, 46]. Related links between entropy production and entanglement in electric-field quenches and pulses have also been discussed recently [28, 22]. In this work, we study in full detail the interplay between different input states, such as thermal and single-mode squeezed states, investigating how they affect the entanglement generated in the Schwinger effect, thus providing a holistic analysis of the Schwinger process Furthermore, we apply the fermionic logarithmic negativity using the separability criterion established in Ref. [71]. Beyond strong QED, recent efforts in collider physics [3, 5, 4] and early-universe cosmology [63, 14] have explored how entanglement-based observables can be used to test quantum correlations in fundamental processes. Similar ideas have been recently discussed in the context of gravitational radiation [51, 52].
@@
The outline of the article is as follows. In Section II we briefly review the Schwinger effect for both scalars and fermions following a mode-by-mode approach in terms of the Bogoliubov coefficients. This approach is particularly useful for understanding quantum entanglement. We focus on two cases: i) constant electric field $\vec{E}$, and ii) constant and anti-parallel electric and magnetic fields $\vec{E}$ and $\vec{B}$. We then introduce the tools needed to quantify quantum entanglement, including the Gaussian formalism, and the definition of logarithmic negativity as our main measure of entanglement in Section III. In Section IV we present our main results: a comparative study of scalar and spinor entanglement in QED, the impact of thermal initial states, and the identification of critical temperatures and electric field strengths above (or under) which entanglement vanishes (for bosons) or is strongly suppressed (for fermions). We also discuss how these findings can be used in future experiments. Finally we summarize our conclusions in Section V.
@@ II Schwinger effect
The instability of the quantum vacuum in strong electric fields was first anticipated by Heisenberg and Euler through an effective action approach for constant backgrounds [41] (see Ref. [79] for scalar QED). Schwinger later formalized this result using the modern language of QED by computing the imaginary part of the one-loop effective action to evaluate the vacuum-persistence amplitude [68],
@@
with $V\Delta t$ the spacetime volume, and $e$, $m$, $E$ the particle charge, mass, and the external field strength, respectively. This expression reveals the exponential suppression of the process for subcritical fields, and defines the natural scale $E_{\mathrm{crit}}=\pi m^{2}/e$ as the threshold where non-perturbative pair production becomes significant (see Ref. [23] for an historical perspective and Ref. [66] for a review on both historical and technical aspects).
In what follows we adopt the (equivalent) mode–function viewpoint based on frequency mixing and Bogoliubov transformations. [^3] See Refs. [50, 21, 27, 12, 9, 8] for different aspects of the Schwinger effect within this formalism, and Ref. [78] for a stochastic generalization in the effective action formalism. See also Ref. [75] for an exact WKB treatment of time-dependent electric fields. For a detailed discussion of quantum field states and their statistical properties in unstable-vacuum settings and critical potential steps, see Refs. [31, 32, 33]. By analyzing the frequency mixing of solutions to the Klein–Gordon and Dirac equations in time-dependent backgrounds, we can identify the natural in and out mode bases and compute the coefficients that relate them. This provides direct access to the particle content of the theory and, crucially for our purposes, to the quantum correlations between the produced pairs.
@@ II.1 Bosons
The appropriate inner product in the space of solutions between two field configurations $\phi_{1}$ and $\phi_{2}$ is given by the conserved symplectic form [^5] The complex-conjugate field carries opposite charge and its (natural) KG product uses $D_{\mu}^{*}=\partial_{\mu}-ieA_{\mu}$. As a result, if $\phi$ has positive KG norm, $\phi^{*}$ has negative KG norm with respect to the conjugate product (see Ref. [10] for further details).
@@ II.2 Fermions
Unlike the bosonic case, and for the sake of brevity, we do not treat here the $\vec{B}=0$ configuration separately, since the Bogoliubov coefficients can be obtained from the case with anti-parallel $\vec{E}$ and $\vec{B}$ fields in the limit $B\to 0$. More details can be found in Appendix A. We proceed by decomposing the spinor field into Fourier modes along the $y$ and $z$ directions, and expanding its $x$-dependence in terms of Hermite polynomials: [^6] The parameter $l$ of the Hermite polynomials is often called “Landau level” (for instance in [21]).
@@ II.3 Bogoliubov coefficients
are the well-known Bogoliubov coefficients, that can be computed as inner products of the in and out solutions[56]. [^7] See Ref. [54] for a detailed computation of the Bogoliubov coefficients in the constant $E$-background case, and for interesting remarks on different gauge choices, the relationship with cosmological particle production, and connection to effective action methods. Using the fact that $|\Gamma(\frac{1}{2}+\frac{i\lambda}{2})|^{2}=\frac{\pi}{\cosh(\frac{\lambda\pi}{2})}e^{\lambda\pi/2}$ one can verify that
@@
Let us move now to the spinor case. Just as in the scalar case, the relationship between the in the out basis comes from the relationship between the different $S_{i}$ functions. In the $\vec{B}\parallel\vec{E}$ case, we find the following Bogoliubov transformations between the creation and annihilation operators at early and late times (in agreement with Refs. [53, 73])
@@
where the Bogoliubov coefficients read (see, for example, Ref. [73] for a detailed derivation)
@@ III Quantum entanglement
As we are interested in quantifying the amount of entanglement in the Schwinger effect, both for bosons and fermions, we provide a pedagogical presentation of the tools and formalism we leverage from quantum information theory. We begin by defining some of the most popular entanglement and correlations quantifiers we use in our calculations, more specifically, entanglement entropy, logarithmic negativity, and mutual information, and then we illustrate how to compute the entanglement in some specific cases of interest.
@@ III.1 Quantifying entanglement
Quantifying the amount of entanglement in quantum states is a topic of central interest in quantum information theory and there exists a whole program of entanglement quantification with several measures [61] being employed for this task. The most widely known, among them, is the von Neumann entropy of the reduced state, computed by $S_{V}=-\text{Tr}_{A}[\hat{\rho}\log_{2}\hat{\rho}$], where $\text{Tr}_{A}$ denotes the partial trace with respect to subsystem $A$ (the result is independent on whether tracing out subsystem $A$ or $B$). This measure, however, is limited to the cases where the total state $\hat{\rho}$ is pure. This is rarely the case in real-life setups since pure states are associated to isolated systems. In realistic frameworks, systems are always exposed, to some degree, to noise resulting from interactions between the system of interest and its environment, i.e. the set of undesired degrees of freedom.
@@
This PPT prescription does not directly apply to fermionic systems, where the associated operators satisfy anticommutation relations rather than commutation relations. In particular, it was shown [71] that using logarithmic negativity for fermionic quantum states based on the PPT criterion discussed in the previous paragraph, which we may also refer to as bosonic PPT, one fails to capture some of the fermionic quantum correlations. The criterion was extended for fermionic systems in [71], where the fermionic partial transposition was introduced, and according to which the fermionic logarithmic negativity was defined to quantify the amount of entanglement in fermionic states. Furthermore, it was shown in [70] that it is an entanglement monotone under local operations and classical communication. The result of the fermionic partial transposition compared to the bosonic one is the introduction of phases in some of the density matrix elements (see section IV.2.3 for the details and Ref. [30] for an example in 1+1.). Finally, to compute entanglement in the fermionic system, we make use of (34) but applied to the density matrix transformed under the fermionic partial transposition.
@@ III.2 Gaussian formalism of continuous variable systems
Continuous variable systems are those whose observables possess a continuous spectrum, such as the harmonic oscillator. Upon quantization, the states of such systems are elements of an infinite-dimensional Hilbert space, and, hence, computing entanglement in such systems amounts to manipulating infinite-dimensional density matrices. Gaussian systems, however, such as those generated by quadratic Hamiltonians, enjoy a drastic simplification: they are fully characterized by their first and second moments, whose size depends on the number of degrees of freedom. We now briefly review the Gaussian formalism and how to employ it to compute entanglement. Given that we are interested in the mode of pairwise entanglement, we will restrict the discussion to the case of two degrees of freedom. A pedagogical reference for the Gaussian state formalism and entanglement in Gaussian states can be found in [69].
@@
Now, let us turn our attention to the logarithmic negativity. Simon extended the PPT criterion to two-mode continuous variable Gaussian systems [72]. The key step towards computing entanglement is the operation of the partial transposition. In terms of a $4\times 4$ covariance matrix $\bm{\sigma}$, the partial transposition reads [69]
@@ III.3 Example: a two-level bipartite system
$\hat{\rho}^{(\text{out})}=\begin{pmatrix}|\alpha|^{2}&0&0&\alpha^{*}\beta^{*}\\ 0&0&0&0\\ 0&0&0&0\\ \alpha\beta&0&0&|\beta|^{2}\end{pmatrix}\,$, (47)
where $\hat{\rho}^{(\mathrm{out})}$ can be derived from $\hat{\rho}^{(\mathrm{in})}$ through the matrix implementation of eq. (46),
$\hat{\rho}^{(\mathrm{out})}=M\hat{\rho}^{(\mathrm{in})}M^{\dagger}\qquad\text{where}\qquad M=\begin{pmatrix}\alpha^{*}&0&0&\beta^{*}\\ 0&1&0&0\\ 0&0&1&0\\ \beta&0&0&\alpha\end{pmatrix}\,$. (48)
@@
where $\omega_{i}$ is the energy of the $i$-th particle and $f_{i}=\left(e^{\omega_{i}/T}+1\right)^{-1}$ is the mean occupation number of the $i$-th particle. We consider $f_{1}=f_{2}=f$. The total state reads
@@
The final density matrix reads
@@
Since the total state is mixed, we cannot use von Neumann entropy to quantify entanglement. Instead, one can use fermionic logarithmic negativity. We do so directly in section IV.2.3 to compute entanglement in the Schwinger effect.